# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ Abstract
Electron spin-qubits in silicon-germanium (SiGe) heterostructures are a major candidate for the realization of scalable quantum computers. A critical challenge in strained Si/SiGe quantum wells (QWs) is the existence of two nearly degenerate valley states at the conduction band minimum that can lead to leakage of quantum information. To address this issue, various strategies have been explored to enhance the valley splitting (i.e., the energy gap between the two low-energy conduction band minima), such as sharp interfaces, oscillating germanium concentrations in the QW (known as wiggle well) and shear strain engineering. In this work, we develop a comprehensive envelope-function theory augmented by empirical nonlocal pseudopotential theory to incorporate the effects of alloy disorder, strain, and non-trivial resonances arising from interaction between valley states across different Brillouin zones. We apply our model to analyze common epitaxial profiles studied in the literature with a focus on wiggle well type structures and compare our results with previous work. Our framework provides an efficient tool for quantifying the interplay of these effects on the valley splitting, enabling complex epitaxial profile optimization in future work.
@@ I Introduction
One of the key challenges in the design of reliable Si/SiGe qubits is the enhancement of the energy splitting between the two nearly degenerate valley states at the conduction band minimum of a biaxially strained SiGe/Si/SiGe quantum well (QW), see Fig. 1. The energy splitting between these states, called valley splitting, is caused by the coupling of the two valley states by the heterostructure potential [15, 16, 17]. Interface roughness and random alloy disorder in the SiGe barrier [18, 19, 20] cause the valley splitting to fluctuate across the chip with typical values ranging from several tens to hundreds of µeV [21, 22, 23, 24]. These statistical fluctuations of the valley splitting are notoriously difficult to control and inevitably lead to spatial domains with very low splitting, where it becomes comparable with the Zeeman-splitting [25, 26, 23]. As a result, low valley splittings lead to so-called spin-valley hotspots, which are a potential source for spin-dephasing and leakage of quantum information [27, 28]. While these issues might be secondary for stationary qubits, they are particularly critical in spin-qubit shuttles, where the electron is conveyed over micrometer distances across a disordered landscape [29, 30]. For such applications a reliably large valley splitting above 200 µeV would be desirable to avoid spin-valley hotspots [31, 29, 32].
@@
Fig. 1: (a) First Brillouin zone of the face-centered cubic (fcc) lattice. The degeneracy between the six equivalent conduction band minima in Si near the $X$-points is lifted by strain. (b) Energy diagram for the conduction band ground states in Si/SiGe quantum dots. Biaxial strain due to lattice mismatch between Si and SiGe leads to a separation of the two valleys oriented along $[001]$ and $[00\overline{1}]$ from the other four conduction band ground state valleys. The heterostructure potential and alloy disorder finally lift the remaining degeneracies. (c) Cubic unit cell of relaxed bulk Si (diamond crystal) composed of two interpenetrating fcc sub-lattices separated by a non-primitive translation along a quarter of the face diagonal. Atoms of the first and second fcc sub-lattice are shown in red and blue, respectively. (d) Biaxial strain along $[100]$ and $[010]$ yields a tetragonal crystal with reduced symmetry. Similar to the relaxed crystal, the two sub-lattices are interchangeable by a nonsymmorphic screw symmetry [31, 33]. (e) Additional shear strain along $[110]$ further reduces the symmetry to an orthorhombic system. The displacement between the two sub-lattices is controlled by Kleinman’s internal ionic displacement parameter $\zeta$ (blue arrows). The nonsymmorphic screw symmetry, which maps the two sub-lattices onto each other, is broken by the shear strain.
@@
Recently, several heuristic strategies have been proposed to enhance the valley splitting by engineering the Si/SiGe heterostructure [34, 31, 19, 33]. These include sharp interfaces, narrow QWs and QWs with uniform low Ge-concentration [18] and more elaborate epitaxial profiles in the QW such as a Ge-spike [35] or an oscillating Ge concentration, known as the wiggle well [34, 31, 33]. In all these concepts, there is a complex interplay of resonances, strain and disorder effects, which can lead to an enhancement of the mean valley splitting. For practical applications, however, it is important to assess the magnitude of the deterministic enhancement along with the strength of the disorder-induced fluctuations.
@@
The theoretical description of valley splitting is primarily based on effective mass models, partially augmented by empirical tight-binding models [36, 37, 19, 18, 33], pseudopotential theory [38, 31] or density functional theory [16, 17, 39]. Moreover, these effective mass models have also been combined with statistical models to incorporate the effects of alloy disorder. For instance, in Refs. [18, 19, 20] alloy fluctuations have been modeled by sampling of Ge atoms in the primitive unit cells of the crystal from a (scaled) binomial distribution. Alternatively, there are also approaches in which the alloy disorder is already taken into account at the level of the electronic band structure [31]. Furthermore, the coupling between valley states across different Brillouin zones is inconsistently accounted for in the literature. These effects lead to non-trivial resonances, that are addressed in several studies [31, 34, 33], whereas most of the existing literature employs the so-called “$2k_{0}$-theory”, where only the direct interaction of the two valley states within the same Brillouin zone is considered [15, 40, 19, 20, 41, 27]. Finally, only few works explicitly discuss the impact of strain [36, 42, 43, 33]. In particular, merely the recent paper by Woods et al. [33] provides a combined analysis of non-trivial resonances and (shear) strain based on an effective mass theory derived from a tight-binding model. This model, however, disregards disorder-induced enhancements of the valley splitting.
@@
In order to improve the understanding of the complex physics determining the valley splitting in Si/SiGe qubits, it is pertinent to develop a comprehensive theoretical model that combines all of the aforementioned effects – namely strain, random alloy disorder and non-trivial resonances – into a unified framework. This objective has been accomplished in the present paper by means of an effective mass model, which merges several existing concepts consistently in a common framework. We demonstrate that our model faithfully reproduces several known results on valley splitting statistics, interface effects and the shear strain-dependency of the long-period wiggle well, but also extends the state of the art.
@@
The paper is organized as follows: In Sec. II , we provide the theoretical model for the valley-splitting in Si/SiGe QDs which involves a multi-valley coupled envelope equation, a statistical model of the random alloy disorder and the emperical pseudopotential method to account for the electronic band structure, strain and crystal symmetries. We provide expressions for the statistical properties of the valley splitting. Numerical results are described in Sec. III. Major attention is devoted to strain-induced effects, which extend previous findings on the QW interface-width dependency and wiggle well-type heterostructures. Finally, Sec. IV provides a thorough discussion of the results and a comparison with similar models followed by an outlook in Sec. V. Several technical considerations on the derivation of the multi-valley coupled envelope equation model, band structure coefficients and valley splitting statistics can be found in the appendix.
@@ II.1 Coupled Envelope Equations
where $\Psi_{\pm}\left(\mathbf{r}\right)$ denotes the envelope wave functions of the corresponding valley states. Here, the valley splitting corresponds to the energy difference between the first excited state and the ground state. The Hamiltonian in Eq. (1) involves
@@
where $m_{l}$ and $m_{t}$ are the effective mass tensor components at the silicon conduction band minimum and $U\left(\mathbf{r}\right)$ is the total confinement potential. The intervalley coupling is described by
$V_{c}\left(\mathbf{r}\right)=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}\,U\left(\mathbf{r}\right),$ (3)
which involves the plane wave expansion coefficients $c_{\pm}\left(\mathbf{G}\right)=c_{\pm\mathbf{k}_{0}}\left(\mathbf{G}\right)$ of the lattice-periodic part of the Bloch factors at the two valleys $u_{\mathbf{\pm k}_{0}}\left(\mathbf{r}\right)=\sum_{\mathbf{G}}\mathrm{e}^{i\mathbf{G}\cdot\mathbf{r}}c_{\mathbf{\pm k}_{0}}\left(\mathbf{G}\right)$. The band index is suppressed throughout this paper, as we are solely concerned with the (lowest-energy) conduction band. A detailed derivation of the coupled envelope equation model (1) can be found in Appendix A.
@@
$U\left(\mathbf{r}\right)=U_{\mathrm{het}}\left(\mathbf{r}\right)+U_{\mathrm{QD}}\left(x,y\right)+U_{F}\left(z\right),$ (4)
describes the effects of both the epitaxial heterostructure and the electrostatic fields induced by the metal gates at the top surface of the device. The heterostructure potential $U_{\mathrm{het}}\left(\mathbf{r}\right)$ describes the potential induced by the Ge atoms in the SiGe alloy, i.e., a type-II Si/SiGe QW with random alloy disorder, which will be described in more detail in Sec. II.2. Note that in the present model, the effects of Ge atoms are entirely described by the heterostructure potential, whereas the underlying band structure (i.e., effective masses and Bloch factors) are those of pure Si. We assume a harmonic QD confinement potential induced by the gate electrodes
@@ II.2 Heterostructure Potential and Alloy Disorder
The heterostructure potential describes the built-in potential due to the epitaxial profile. In order to account for disorder effects in the Si1-xGex alloy, we choose a statistical model similar to that in Refs. [20, 41, 44], where the heterostructure potential is described as a random field
@@
is a smoothed indicator function that describes the shape of the QW. Here, $h$ is the QW thickness and $\sigma_{u}$ and $\sigma_{l}$ describe the width of the upper and lower QW interfaces, respectively. The interface widths can be obtained experimentally using scanning transmission electron microscopy, typical values for Si/SiGe QWs are $\sigma_{u}\approx\sigma_{l}\approx 0.5\,\mathrm{nm}$ [18, 44, 32].
@@
$\approx\left(\Delta E_{c}\right)^{2}\Omega_{a}\,X_{\mathrm{}}\left(z\right)\left(1-X\left(z\right)\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right).$
The simple covariance function reflects the assumption of locally independent distribution of Ge atoms stated in Eq. (8) (i.e., no clustering of Ge atoms) and determines the statistical properties of the intervalley-coupling parameter, see Appendix C for details.
@@ II.3 Empirical Pseudopotential Theory and Strain
The empirical pseudopotential method (EPM) provides an accurate description of the electronic band structure with only a few parameters fitted to experimental data. The approach is particularly efficient for silicon, germanium and other semiconductors [47, 48, 49]. A major advantage of the EPM is that it can naturally account for strain effects arising from a displacement of the crystal ions $\mathbf{R}_{i}^{\prime}=\left(I+\varepsilon\right)\mathbf{R}_{i}$, where $\varepsilon$ is the strain tensor. The key steps for the inclusion of strain in EPM calculations are the consideration of strained reciprocal lattice vectors $\mathbf{G}_{i}^{\prime}\approx\left(I-\varepsilon\right)\mathbf{G}_{i}$ (assuming small strain to linear order), the modification of the primitive unit cell volume $\Omega_{p}^{\prime}\approx\left(1+\mathrm{tr}\left(\varepsilon\right)\right)\Omega_{p}$, interpolation of the atomic pseudopotentials at the strained lattice vectors and consideration of internal ionic displacement, describing the relative shift between the two sub-lattices, see Fig. 1 (e). For strained SiGe alloys, numerous empirical pseudopotential models can be found in the literature [50, 51, 52, 53, 54, 43, 55].
In this paper, we use the nonlocal EPM model described by Fischetti & Laux [51]. From this model, we have obtained the plane wave expansion coefficients of the Bloch factors $c_{\mathbf{k}}\left(\mathbf{G}\right)$, the location of the conduction band minima ±$\pm\mathbf{k}_{0}=\left(0,0,\pm k_{0}\right)$ and the corresponding effective mass tensor components $m_{l}$ and $m_{t}$ by numerical diagonalization of the pseudopotential Hamiltonian (taking 113 plane waves into account). We assume biaxial (tensile) strain due to the lattice mismatch between the $\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ substrate and the Si quantum well [56]
@@
with $\varepsilon_{\parallel}=\varepsilon_{x,x}=\varepsilon_{y,y}=a_{0}^{\mathrm{SiGe}}/a_{0}^{\mathrm{Si}}-1\approx 1.14\,\%$ and $\varepsilon_{\perp}=\varepsilon_{z,z}=-2C_{12}/C_{11}\varepsilon_{\parallel}\approx-0.88\,\%$, where $C_{1,2}$ and $C_{2,2}$ are elastic constants of Si [51]. Below, we will consider additional shear strain, which might originate from the macroscopic device geometry [57, 58, 33], dislocations [59, 60] or alloy disorder [61, 62, 63].

### v3

@@ Abstract
Electron spin-qubits in silicon-germanium (SiGe) heterostructures are a major candidate for the realization of scalable quantum computers. A critical challenge in strained Si/SiGe quantum wells (QWs) is the existence of two nearly degenerate valley states at the conduction band minimum that can lead to leakage of quantum information. To address this issue, various strategies have been explored to enhance the valley splitting (i.e., the energy gap between the two low-energy conduction band minima), such as sharp interfaces, oscillating germanium concentrations in the QW (known as wiggle wells) and shear strain engineering. In this work, we develop a comprehensive envelope-function theory augmented by an empirical nonlocal pseudopotential model to incorporate the effects of alloy disorder, strain, and non-trivial resonances arising from interactions between valley states across neighboring Brillouin zones. We apply our model to analyze common epitaxial profiles studied in the literature with a focus on wiggle well type structures and compare our results with previous work. Our framework provides an efficient tool for quantifying the interplay of these effects on the valley splitting, enabling complex epitaxial profile optimization in future work.
@@ I Introduction
One of the key challenges in the design of reliable Si/SiGe qubits is the enhancement of the energy splitting between the two nearly degenerate valley states at the conduction band minimum of a biaxially strained SiGe/Si/SiGe quantum well (QW), see Fig. 1. The energy splitting between these states, called valley splitting, is caused by the coupling of the two valley states by the heterostructure potential [15, 16, 17]. Interface roughness and random alloy disorder in the SiGe barrier [18, 19, 20] cause the valley splitting to fluctuate across the chip with typical values ranging from several tens to hundreds of µeV [21, 22, 23, 24]. These statistical fluctuations of the valley splitting are notoriously difficult to control and inevitably lead to spatial domains with very low splitting, where it becomes comparable with the Zeeman-splitting [25, 26, 23]. As a result, low valley splittings lead to so-called spin-valley hotspots, which are a potential source for spin-dephasing and leakage of quantum information [27, 28]. While these issues might be secondary for stationary qubits, they are particularly critical in spin-qubit shuttles, where the electron is conveyed over micrometer distances across a disordered landscape [29, 30]. For such applications a reliably large valley splitting is desirable to avoid spin-valley hotspots [31, 29, 32].
@@
Fig. 1: (a) First Brillouin zone of the face-centered cubic (fcc) lattice. The degeneracy between the six equivalent conduction band minima in Si near the $X$-points is lifted by strain. (b) Energy diagram for the conduction band ground states in Si/SiGe quantum dots. Biaxial strain due to lattice mismatch between Si and SiGe leads to a separation of the two valleys oriented along [001][001][ 001 ] and $[00\overline{1}]$ from the other four conduction band ground state valleys. The heterostructure potential and alloy disorder finally lift the remaining degeneracies. (c) Cubic unit cell of relaxed bulk Si (diamond crystal) composed of two interpenetrating fcc sub-lattices separated by a non-primitive translation along a quarter of the face diagonal. Atoms of the first and second fcc sub-lattice are shown in red and blue, respectively. (d) Biaxial strain along [100][100][ 100 ] and [010][010][ 010 ] yields a tetragonal crystal with reduced symmetry. Similar to the relaxed crystal, the two sub-lattices are interchangeable by a nonsymmorphic screw symmetry [31, 33]. (e) Additional shear strain along [110][110][ 110 ] further reduces the symmetry to an orthorhombic system. The displacement between the two sub-lattices is controlled by Kleinman’s internal ionic displacement parameter $\zeta$ (blue arrows). The nonsymmorphic screw symmetry, which maps the two sub-lattices onto each other, is broken by the shear strain.
@@
Recently, several heuristic strategies have been proposed to enhance the valley splitting by engineering the Si/SiGe heterostructure [34, 31, 19, 33]. These include sharp interfaces, narrow QWs and QWs with uniform low Ge-concentration [18] and more advanced epitaxial profiles in the QW such as a Ge-spike [35] or an oscillating Ge concentration, known as the wiggle well [34, 31, 33, 36]. In all these concepts, there is a complex interplay of resonances, strain and disorder effects, which can lead to an enhancement of the mean valley splitting. For practical applications, however, it is important to assess the magnitude of the deterministic enhancement along with the strength of the disorder-induced fluctuations.
@@
In the literature, the theoretical description of valley splitting is primarily based on envelope function theory and effective mass type models, partially augmented by empirical tight-binding models [37, 38, 19, 18, 33], pseudopotential theory [39, 31] or density functional theory [16, 17, 40]. Moreover, these models have been combined with statistical models to incorporate the effects of alloy disorder. For instance, in Refs. [18, 19, 20] alloy fluctuations have been modeled by sampling of Ge atoms in the primitive unit cells of the crystal from a (scaled) binomial distribution. Alternatively, there are also approaches in which the alloy disorder is already taken into account at the level of the electronic band structure [31]. Furthermore, the coupling between valley states across different Brillouin zones is inconsistently accounted for in the literature. These effects lead to non-trivial resonances, that are addressed in several studies [31, 34, 33], whereas most of the existing literature employs the so-called “$2k_{0}$-theory”, where only the direct interaction of the two valley states within the same Brillouin zone is taken into account [15, 41, 19, 20, 42, 27]. Finally, only few works explicitly discuss the impact of strain [37, 43, 44, 33]. In particular, merely the recent paper by Woods et al. [33] provides a combined analysis of non-trivial resonances and (shear) strain based on an effective mass theory derived from a tight-binding model. Their work, however, disregards the effects of alloy disorder.
@@
In order to improve the understanding of the complex physics determining the valley splitting in Si/SiGe qubits, it is pertinent to develop a comprehensive theoretical model that combines all of the aforementioned effects – namely strain, random alloy disorder and non-trivial resonances – into a unified framework. This objective has been accomplished in the present work by means of an envelope function model, which merges several existing concepts consistently in a common framework. We demonstrate that our model faithfully reproduces several known results on valley splitting statistics, interface effects and the shear strain-dependency of the long-period wiggle well, but also extends the state of the art.
@@
The paper is organized as follows: In Sec. II , we provide the theoretical model for the valley-splitting in Si/SiGe QDs which involves a multi-valley coupled envelope equation, a statistical model of the random alloy disorder and the empirical pseudopotential method to account for the electronic band structure, strain and crystal symmetries. We provide expressions for the statistical properties of the valley splitting. Numerical results are described in Sec. III. Major attention is devoted to strain-induced effects, which extend previous findings on the QW interface-width dependency and wiggle well-type heterostructures. Finally, Sec. IV provides a thorough discussion of the results and a comparison with similar models followed by an outlook in Sec. V. Several technical considerations on the derivation of the multi-valley coupled envelope equation model, the pseudopotential model, band structure coefficients and statistics can be found in the appendix.
@@ II.1 Coupled Envelope Equations
where $\Psi_{\pm}\left(\mathbf{r}\right)$ denotes the envelope wave functions of the corresponding valley states. Here, the valley splitting corresponds to the energy difference between the first excited state and the ground state. The Hamiltonian takes the form
@@
where $m_{l}$ and $m_{t}$ are the effective mass tensor components at the silicon conduction band minimum and $U\left(\mathbf{r}\right)$ is the total confinement potential. Since the effective mass parameters of conduction band electrons in Si1-xGex alloys are practically constant for Ge content $x\apprle 0.85$ [45], we assume the same effective masses in the QW and in the barrier. The intervalley coupling is described by
$V_{c}\left(\mathbf{r}\right)$ $=\mathrm{e}^{-2i\mathbf{k}_{0}\cdot\mathbf{r}}u_{+}^{*}\left(\mathbf{r}\right)u_{-}\left(\mathbf{r}\right)U\left(\mathbf{r}\right)$ (3)
$=\sum_{\mathbf{G},\mathbf{G}^{\prime}}\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)U\left(\mathbf{r}\right),$
which involves the plane wave expansion coefficients $c_{\pm}\left(\mathbf{G}\right)=c_{\pm\mathbf{k}_{0}}\left(\mathbf{G}\right)$ of the lattice-periodic part of the Bloch factors $u_{\pm}\left(\mathbf{r}\right)=\sum_{\mathbf{G}}\mathrm{e}^{i\mathbf{G}\cdot\mathbf{r}}c_{\pm}\left(\mathbf{G}\right)$ at the two valleys. The band index is suppressed throughout this paper, as we are solely concerned with the (lowest energy) conduction band. A detailed derivation of the coupled envelope equation model (1) is given in Appendix A.
@@
$U\left(\mathbf{r}\right)=U_{\mathrm{het}}\left(\mathbf{r}\right)+U_{\mathrm{QD}}\left(x,y\right)+U_{F}\left(z\right)$ (4)
describes the effects of both the epitaxial heterostructure and the electrostatic fields induced by the metal gates at the top surface of the device. The heterostructure potential $U_{\mathrm{het}}\left(\mathbf{r}\right)$ models the potential induced by the Ge atoms in the SiGe alloy, i.e., a type-II Si/SiGe QW with random alloy disorder, which will be described in more detail in Sec. II.2 below. Note that in the present model, the effects of Ge atoms are entirely described by the heterostructure potential, whereas the underlying band structure coefficients (i.e., effective masses and Bloch factors) are those of pure Si  [^1] In principle, one could also consider a (small) non-zero Ge concentration in the bulk potential $V\left(\mathbf{r}\right)$, see Eq. (40), and treat the resulting Si1-xGex alloy using the virtual crystal approximation (VCA). For consistency, the same Ge concentration must then be removed from the mean of the heterostructure potential $U\left(\mathbf{r}\right)$. While this approach could yield some enhancements for the band structure coefficients and effective masses, the VCA will qualitatively preserve the symmetries of the diamond crystal [31], leading to a similar shear strain dependency. For low Ge concentrations, only minor changes in the quantitative results are expected since the effective mass and band structure parameters of the Si1-xGex-alloy (in VCA) are very close to the pure Si values [54, 45].. We assume a harmonic QD confinement potential induced by the gate electrodes
@@ II.2 Heterostructure Potential and Alloy Disorder
The heterostructure potential describes the built-in potential due to the epitaxial profile. In order to account for disorder in the Si1-xGex alloy, we choose a statistical model similar to that in Refs. [20, 42, 47], where the heterostructure potential is described as a random field
@@
is a smoothed indicator function that describes the shape of the QW. Here, $h$ is the QW thickness and $\sigma_{u}$ and $\sigma_{l}$ describe the width of the upper and lower QW interfaces, respectively. The interface widths can be obtained experimentally using scanning transmission electron microscopy. Typical values for Si/SiGe QWs are $\sigma_{u}\approx\sigma_{l}\approx 0.5\,\mathrm{nm}$ [18, 47, 32].
@@
$\approx\left(\Delta E_{c}\right)^{2}\Omega_{a}\,X\left(z\right)\left(1-X\left(z\right)\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right).$
The covariance function reflects the assumption of locally independent distribution of Ge atoms stated in Eq. (8) (i.e., no clustering of Ge atoms) and determines the statistical properties of the intervalley-coupling parameter, see Appendix D for details.
@@ II.3 Empirical Pseudopotential Theory and Strain
The empirical pseudopotential method (EPM) provides an accurate description of the electronic band structure with only a few parameters fitted to experimental data [49, 50, 51]. A major advantage of the EPM is that it can naturally account for strain effects arising from a displacement of the crystal ions $\mathbf{R}_{i}^{\prime}=\left(I+\varepsilon\right)\mathbf{R}_{i}$, where $\varepsilon$ is the strain tensor. For strained SiGe alloys, numerous empirical pseudopotential models are available in the literature [52, 53, 54, 55, 56, 44, 57]. The key steps for the inclusion of strain in EPMs are [55]:
  1. 1.
    computation of strained reciprocal lattice vectors $\mathbf{G}_{i}^{\prime}\approx\left(I-\varepsilon\right)\mathbf{G}_{i}$ (assuming small strain to linear order), where $\textbf{G}_{i}$ denotes the reciprocal lattice vectors of the relaxed crystal
  2. 2.
    strain-induced modification of the primitive unit cell volume $\Omega_{p}^{\prime}\approx\left(1+\mathrm{tr}\left(\varepsilon\right)\right)\Omega_{p}$
  3. 3.
    interpolation of the Fourier coefficients of the pseudopotential at strained reciprocal lattice vectors
  4. 4.
    consideration of internal ionic displacement, see Fig. 1 (e).
In this paper, we employ the nonlocal empirical pseudopotential model described by Rieger & Vogl [54]. From this model, we have obtained the plane wave expansion coefficients of the Bloch factors $c_{\mathbf{k}}\left(\mathbf{G}\right)$, the wave numbers of the conduction band minima ±$\pm\mathbf{k}_{0}=\left(0,0,\pm k_{0}\right)$ and the corresponding effective mass tensor components $m_{l}$ and $m_{t}$, see Tab. 1. We have used 181 plane waves corresponding to a cutoff energy of $12\,\mathrm{Ry}$ for convergence [54]. Details on the EPM are given in Appendix B.
The theory presented in this paper is limited to spatially homogeneous strain distributions. If the strain field is only slowly varying (on a length scale that is large compared to the characteristic size of the QD envelope wave function), the model can be applied locally as is. For spatially rapidly varying strain fields, a generalization of the model might be required. We assume biaxial (tensile) strain due to the lattice mismatch between the $\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ substrate and the Si QW [58]
@@
with $\varepsilon_{\parallel}=\varepsilon_{x,x}=\varepsilon_{y,y}=a_{0}^{\mathrm{SiGe}}/a_{0}^{\mathrm{Si}}-1\approx 1.14\,\%$ and $\varepsilon_{\perp}=\varepsilon_{z,z}=-2C_{1,2}/C_{1,1}\varepsilon_{\parallel}\approx-0.88\,\%$, where $C_{1,1}$ and $C_{1,2}$ are elastic constants of Si, see Tab. B. Below, we will consider additional shear strain along the [110] crystallographic direction. Our analysis of shear strain induced effects will be purely phenomenological, i.e., we will not discuss how it might be generated. We remark that the shear strain can be engineered by design of the macroscopic device geometry [44, 59, 60, 33], but it might also originate from crystal defects [61, 62] or alloy disorder [63, 64]. Effective masses and the conduction band minimum wave numbers are evaluated specifically for the strain tensors considered in the simulations.
symbol description value
$k_{0}$ conduction band min. wave number $0.839\times 2\pi/a_{0}$
$m_{t}$ transverse effective mass $0.202\times m_{0}$
$m_{l}$ longitudinal effective mass $0.920\times m_{0}$
Tab. 1: Band parameters for the conduction band minimum computed from the pseudopotential model for relaxed bulk Si. The wave number $k_{0}$ has been obtained from minimization of the conduction band energy. The effective masses were computed using a finite difference approximation of the band curvature at $\mathbf{k}_{0}$. Parameter values used in the simulations were specifically evaluated for the respective strain tensors.